# Moving Mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) Method for Time-dependent Convection-Dominated Convection-Diffusion Problems

### v3

@@ 3 Results for Isotropic Diffusion
$u_{0}(x,y)=\textrm{e}^{-\frac{(x-1)^{2}+y^{2}}{0.25}}+\textrm{e}^{-\frac{(x-1)^{2}+(y-0.5)^{2}}{0.25}}+\textrm{e}^{-\frac{(x-1)^{2}+(y+0.5)^{2}}{0.25}},$ (35)
@@
The results are shown in Figure 1 with side view profiles and Figure 2 with top view profiles. From Figures 1(a) and (c), we see that for the classic Galerkin method, both FM-FEM and MM-FEM, the numerical solutions have non-physical oscillations form near the outflow boundary at $x=10$. Figure 1(b) shows that FM-SUPG resolves the boundary layers present in FM-FEM and MM-FEM but has artificial oscillations on the interior of $\Omega$. On the other hand, Figure 1(d) shows that MM-SUPG results in the smoothest surface among all four methods, and both interior oscillations and oscillations near the outflow have been resolved. Even with smaller $\Delta t$ as $5\times 10^{-3}$, we see in Figure 2 that MM-FEM still has layers at the outflow, where MM-SUPG performs much better.
@@
×$\times\left(\frac{1}{2}+\frac{1}{\pi}\arctan\left(2\varepsilon^{-1/2}(0.25^{2}-(x-0.5)^{2}-(y-0.5)^{2})\right)\right),$ (38)
@@
For coarse mesh with fewer elements (that is, smaller $N$), we observe artificial oscillations in the direction of ${\boldsymbol{b}}$, including MM-FEM, for example, with $N=8192$. Increasing the number of mesh elements helps to improve the results for MM-FEM. Figure 3 shows the top view of the solutions at $T=0.5$ for the four different methods with $N=32768$, while Figure 4 displays the side view of the solution profiles. As can be seen, significant smearing occurs in the solution obtained using the FM-FEM and FM-SUPG, though the latter shows an improvement. On the other hand, both the MM-FEM and MM-SUPG perform well with $N=32768$ elements in the mesh. This is supported by the $H^{1}$-seminorm errors in Table 1 and cap view in Figure 5. The $L^{2}$-norm errors in Table 2 shows that MM-SUPG performs a bit better than MM-FEM for this example. We observe pronounced mesh movement in the direction of ${\boldsymbol{b}}$ for MM-FEM and MM-SUPG, as shown in Figure 6. The rate of convergence in the $H^{1}$-seminorm and $L^{2}$ norm are shown in Figure 7.
@@
Figure 3: Example 3.2 – top view with four different methods with $\Delta t=0.001$ for ${\boldsymbol{b}}=(2,3)$, $N=32768$ at $T=0.5$.
@@
Figure 5: Example 3.2 – cap view moving mesh with $\Delta t=0.001$, ${\boldsymbol{b}}=(2,3)$, $N=32768$ at $T=0.5$.
Refer to caption (a): MM-FEM           Refer to caption (b): MM-SUPG
@@
131072 0.0004 0.0003 0.1995 0.0180
Within this example, we also consider a time-dependent flow given by ${\boldsymbol{b}}=(y-t,x-t)$. The convergence rates in the $H^{1}$-seminorm and $L^{2}$ norm can be found in Figure 8, where numerical values are recorded in Tables 3 and 4. Again, both MM-FEM and MM-SUPG perform much better than FM-FEM and FM-SUPG with larger $N$.
@@
We see small oscillations along the top of the surface in the direction of the outflow in Figures 5 and 9, even for large $N$. However, for larger $N$, we see that DMP is not violated. Both methods converge to the same approximate solution as $N$ increases. In Figure 4 we see that the MM-SUPG method is a noticeable improvement compared to all methods. In Figure 9, only slight oscillations are observable on the backside of the surface (in the direction of the streamlines) for MM-FEM. This is more prominent for smaller $N$ in the MM-FEM case. Overall, MM-SUPG performs the best among all the four methods under consideration.
@@ 4 Discrete Maximum Principle for Anisotropic Diffusion
where ${\boldsymbol{n}}_{i}^{K}$ is the unit inward normal pointing to ${\boldsymbol{a}}_{i}^{K}$, and $h_{i}^{K}$ is the distance from vertex ${\boldsymbol{a}}_{i}$ to surface $S_{i}$. Figure 10 shows an illustration of ${\boldsymbol{q}}$ vectors.
@@
We also have the following inequality due to the Divergence Theorem:
$\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}=-\frac{1}{2}\int_{K}(\nabla\cdot{\boldsymbol{b}})\varphi_{i}^{2}\textrm{d}{\boldsymbol{x}}.$ (43)
Following [26], we define the diffusion-dependent map $\mathbb{D}_{K}^{-1/2}{\boldsymbol{x}}:K\to\tilde{K}$. Define the $\mathbb{D}_{K}$-norm by $\|{\boldsymbol{x}}\|_{\mathbb{D}_{K}}=\sqrt{{\boldsymbol{x}}^{T}\mathbb{D}_{K}{\boldsymbol{x}}}$. Then we have the following relations
@@
Denote by $\tilde{\alpha}_{ij}^{K}$ the dihedral angle between surfaces $\tilde{S}_{i}^{K}$ and $\tilde{S}_{j}^{K}$. Then
@@
The difficulty with the finite element formulation mentioned in Section 2 is the coupling of the diffusion tensor with the flow field ${\boldsymbol{b}}$ in (27) as follows:
@@
where $\mathbb{D}_{i}$ is the vector $i$th row of $\mathbb{D}$. An essential ingredient to the following analysis is that the tensor defined by
$\mathbb{K}=\sum_{i=1}^{d}\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}\otimes\mathbb{D}_{(i)}$ (50)
is symmetric and positive definite. This highlights a significant relationship between $\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}$ and the $i$th row of $\mathbb{D}$ when using SUPG. The added diffusion integral term in (6) can be written in general form as
$\tau_{K}\int_{K}(\mathbb{D}\nabla u)\cdot\nabla({\boldsymbol{b}}\cdot\nabla v)\textrm{d}{\boldsymbol{x}}$ $=\tau_{K}\sum_{i=1}^{d}\int_{K}(\nabla v)^{T}\left(\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}\otimes\mathbb{D}_{(i)}\right)\nabla u\textrm{d}{\boldsymbol{x}}$ (51)
One may view $\mathbb{K}$ as a kind of weighted diffusion tensor.
@@
$(\nabla\varphi_{i})^{T}\mathbb{F}_{K}\nabla\varphi_{j}=\begin{cases}\displaystyle-\frac{\cos(\tilde{\alpha}_{ij}^{K})}{\tilde{h}_{i}^{K}\tilde{h}_{j}^{K}}\textrm{\ \ for \ \ }i\neq j\\ \displaystyle\frac{1}{(\tilde{h}_{i}^{K})^{2}}\textrm{\ \ for \ \ i = j}.\end{cases}$ (54)
@@ 4.1 General flow forms
$\frac{\partial{\boldsymbol{b}}}{\partial x_{1}}\otimes\mathbb{D}_{(1)}=\begin{pmatrix}\displaystyle\frac{\partial b_{1}}{\partial x_{1}}D_{11}&\displaystyle\frac{\partial b_{1}}{\partial x_{1}}D_{12}\\ \displaystyle\frac{\partial b_{2}}{\partial x_{1}}D_{11}&\displaystyle\frac{\partial b_{2}}{\partial x_{1}}D_{12}\end{pmatrix}$ (56)
@@
$\frac{\partial b_{1}}{\partial x_{1}}$ $=\left(\frac{D_{11}}{D_{12}}\right)\frac{\partial b_{2}}{\partial x_{1}}$ (64)
@@
$\frac{\partial b_{1}}{\partial x_{2}}$ $=\left(\frac{D_{21}}{D_{22}}\right)\frac{\partial b_{2}}{\partial x_{2}}$ (66)
@@
$\left(\frac{D_{11}}{D_{12}}\right)b_{2}(x_{1},x_{2},t)+f_{2}(x_{2},t)$ $=\left(\frac{D_{21}}{D_{22}}\right)b_{2}(x_{1},x_{2},t)+f_{1}(x_{1},t)$ (68)
$\left(\frac{D_{11}}{D_{12}}-\frac{D_{21}}{D_{22}}\right)b_{2}(x_{1},x_{2},t)$ $=f_{1}(x_{2},t)-f_{2}(x_{1},t)$ (69)
@@
$b_{1}(x_{1},x_{2},t)$ $=\left(\frac{D_{11}}{D_{12}}\right)\frac{D_{12}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))}{\det(\mathbb{D})}+f_{2}(x_{2},t)$ (71)
$=\frac{D_{11}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))+f_{2}(x_{2},t)(D_{11}D_{22}-D_{12}D_{21})}{\det(\mathbb{D})}$ (72)
@@
$\frac{1}{\det(\mathbb{D})}\left(D_{22}f_{1}(x_{1},t)\mathbb{D}_{(1)}-D_{12}f_{2}(x_{2},t)\mathbb{D}_{(2)}\right)$ (74)
@@
$\frac{1}{\det(\mathbb{D})}\left(D_{11}g_{2}(x_{2},t)\mathbb{D}_{(2)}-D_{21}g_{1}(x_{1},t)\mathbb{D}_{(1)}\right)$ (75)
@@
${\boldsymbol{b}}(x_{1},x_{2},t)=\begin{pmatrix}\displaystyle\frac{f(x_{1},t)}{D_{11}}\\ \displaystyle\frac{g(x_{2},t)}{D_{22}}\end{pmatrix}$ (76)
@@
$\frac{D_{22}}{\det(\mathbb{D})}\left(D_{11}\frac{\partial f_{1}}{\partial x_{1}}-D_{12}\frac{\partial f_{2}}{\partial x_{2}}\right).$ (78)

### v4

@@ 3 Results for Isotropic Diffusion
Note that for $\varepsilon=O(1)$, all methods produce similar results, confirming that stabilization and mesh adaptation are primarily beneficial in convection-dominated regimes.
@@
$u_{0}(x,y)=\textrm{e}^{-\frac{(x-1)^{2}+y^{2}}{0.25}}+\textrm{e}^{-\frac{(x-1)^{2}+(y-0.5)^{2}}{0.25}}+\textrm{e}^{-\frac{(x-1)^{2}+(y+0.5)^{2}}{0.25}}.$ (36)
@@
The results are shown in Figure 1 with side view profiles. From Figures 1(a) and (c), we see that for the classic Galerkin method, both FM-FEM and MM-FEM, the numerical solutions have non-physical oscillations form near the outflow boundary at $x=10$. Figure 1(b) shows that FM-SUPG resolves the boundary layers present in FM-FEM and MM-FEM but has artificial oscillations on the interior of $\Omega$. On the other hand, Figure 1(d) shows that MM-SUPG results in the smoothest surface among all four methods, and both interior oscillations and oscillations near the outflow have been resolved. The computations are also performed with smaller $\Delta t$ as $5\times 10^{-3}$, and the results are shown in Figure 2. As can be seen, even with smaller time step, MM-FEM still has layers at the outflow while MM-SUPG performs much better.
@@
$\times\left(\frac{1}{2}+\frac{1}{\pi}\arctan\left(2\varepsilon^{-1/2}(0.25^{2}-(x-0.5)^{2}-(y-0.5)^{2})\right)\right),$ (39)
@@
For coarse mesh with fewer elements (that is, smaller $N$), we observe artificial oscillations in the direction of $\boldsymbol{b}$}, including MM-FEM, for example, with $N=8192$. Increasing the number of mesh elements helps to improve the results for MM-FEM. Figure 3 shows the top view of the solutions at $T=0.5$ for the four different methods with $N=32768$, while Figure 4 displays the side view of the solution profiles. As can be seen, significant smearing occurs in the solution obtained using the FM-FEM and FM-SUPG, though the latter shows an improvement. On the other hand, both the MM-FEM and MM-SUPG perform well with $N=32768$ elements in the mesh. This is supported by the $H^{1}$-seminorm errors in Table 1 and numerical solutions shown in Figure 4. The $L^{2}$-norm errors in Table 2 shows that MM-SUPG performs comparable even better than MM-FEM for this example. We observe pronounced mesh movement in the direction of $\boldsymbol{b}$} for MM-FEM and MM-SUPG, as shown in Figure 5. The rate of convergence in the $H^{1}$-seminorm and $L^{2}$ norm, based on errors from Tables 1 and 2 are shown in Figure 6. The reference slopes correspond to expected convergence rates: $O(h)$ in $H^{1}$ and $O(h^{3/2}$ in $L^{2}$ using MMPDE framework for convection-dominated problems.
@@
Figure 3: Example 3.2 – top view with four different methods with $\Delta t=0.001$ for ${\boldsymbol{b}}=(2,3)$, $N=32768$ at $T=0.5$. The observed discreteness in (a) is due to the use of piecewise linear finite elements on a fixed mesh where numerical oscillation is significant, which results into obvious color changes in the top view of the solution.
@@
Figure 6: Example 3.2 – Errors corresponds to Tables 1 and 2. $H^{1}$-seminorm and $L^{2}$ norm errors with $\varepsilon=10^{-6}$ and $\Delta t=0.001$ with flow field ${\boldsymbol{b}}=(2,3)$ at $T=0.5$.
@@
131072 0.0004 0.0005 0.1995 0.0180
Within this example, we also consider a time-dependent flow given by ${\boldsymbol{b}}=(y-t,x-t)$. This time-dependent velocity field introduces rotational and time-varying transport, testing robustness of the method under non-uniform flow. The convergence rates in the $H^{1}$-seminorm and $L^{2}$ norm can be found in Figure 7, where numerical values are recorded in Tables 3 and 4. Again, both MM-FEM and MM-SUPG perform much better than FM-FEM and FM-SUPG with larger $N$.
@@
Figure 7: Example 3.2 – Errors corresponds to Tables 3 and 4. $H^{1}$-seminorm and $L^{2}$ norm errors with $\varepsilon=10^{-6}$ and $\Delta t=0.001$ with flow field ${\boldsymbol{b}}=(y-t,x-t)$ at $T=0.5$.
@@
We see small oscillations along the top of the surface in the direction of the outflow in Figures 8, even for large $N$. However, for larger $N$, we see that DMP is not violated. Both methods converge to the same approximate solution as $N$ increases. In Figure 4 we see that the MM-SUPG method is a noticeable improvement compared to all methods. In Figure 8, only slight oscillations are observable on the backside of the surface (in the direction of the streamlines) for MM-FEM. This is more prominent for smaller $N$ in the MM-FEM case. The different results are due to the fact that Example 3.1 involves strong advective transport of sharp features, where SUPG stabilization plays a dominant role. In contrast, Example 3.2 has a smoother structure where mesh adaptation alone already captures the solution well, reducing the relative impact of SUPG. Overall, MM-SUPG provides improved stability and reduced oscillations compared to fixed mesh methods, while offering comparable accuracy to MM-FEM.
@@ 4 Discrete Maximum Principle for Anisotropic Diffusion
In the isotropic case, mesh adaptation primarily improves resolution of steep gradients. In contrast, for anisotropic diffusion, the orientation of the mesh relative to the diffusion tensor becomes critical. The discrete maximum principle is highly sensitive to mesh geometry and coefficient alignment. In the MM-SUPG framework, both the stiffness matrix and stabilization terms depend on the evolving mesh. Misalignment may lead to violation of the discrete maximum principle (DMP), even in stabilized formulations. Therefore, the design of metric tensors must account for both diffusion anisotropy and convection direction. The proposed tensor and metric intersection strategy aim to balance the competing effects of convection and anisotropic diffusion.
@@
where ${\boldsymbol{n}}_{i}^{K}$ is the unit inward normal pointing to ${\boldsymbol{a}}_{i}^{K}$, and $h_{i}^{K}$ is the distance from vertex ${\boldsymbol{a}}_{i}^{K}$ to face $S_{i}^{K}$. Figure 9 shows an illustration of $\boldsymbol{q}$} vectors.
@@
Using the Divergence Theorem on an element $K$, we have
$\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}=\frac{1}{2}\int_{\partial K}\varphi_{i}^{2}({\boldsymbol{b}}\cdot{\boldsymbol{n}})\textrm{d}{\boldsymbol{s}}-\frac{1}{2}\int_{K}(\nabla\cdot{\boldsymbol{b}})\varphi_{i}^{2}\textrm{d}{\boldsymbol{x}}.$ (44)
Equation (44) does not determine the sign of the local convection term in general, since the boundary flux contribution on $\partial K$ need not vanish elementwise. Therefore, in the following analysis, we do not use (44) to infer positivity or negativity of $\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}$. Instead, we use quantitative bounds as follows. Using Cauchy–Schwarz inequality, we obtain
$\left|\int_{K}\phi_{j}(b\cdot\nabla\phi_{i})\,dx\right|$ $\leq\|\phi_{j}\|_{L^{2}(K)}\|b\cdot\nabla\phi_{i}\|_{L^{2}(K)}$ (45)
$\leq\|\phi_{j}\|_{L^{2}(K)}\|b\|_{\infty,K}\|\nabla\phi_{i}\|_{L^{2}(K)}.$ (46)
Since
$\int_{K}\phi_{j}^{2}\,dx=\frac{2|K|}{(d+1)(d+2)}\quad\text{and}\quad\|\nabla\phi_{i}\|_{L^{2}(K)}^{2}=\frac{|K|}{(h_{i}^{K})^{2}},$
it follows that
$\left|\int_{K}\phi_{j}(b\cdot\nabla\phi_{i})\,dx\right|\leq\frac{\sqrt{2}\,|K|}{\sqrt{(d+1)(d+2)}\,h_{i}^{K}}\,\|b\|_{\infty,K}.$ (47)
A simpler estimate is obtained by noting that $\nabla\phi_{i}$ is constant on $K$, $\int_{K}\varphi_{i}\textrm{d}{\boldsymbol{x}}=\frac{K}{d+1}$, and $|\nabla\varphi_{i}|=\frac{1}{h_{i}^{K}}$:
$\left|\int_{K}\phi_{j}(b\cdot\nabla\phi_{i})\,dx\right|$ $\leq\int_{K}|\phi_{j}|\,|b|\,|\nabla\phi_{i}|\,dx$ (48)
$\leq\|b\|_{\infty,K}|\nabla\phi_{i}|\int_{K}\phi_{j}\,dx=\frac{|K|}{(d+1)h_{i}^{K}}\,\|b\|_{\infty,K}.$
Thus,
$\left|\int_{K}\phi_{j}(b\cdot\nabla\phi_{i})\,dx\right|\leq\frac{|K|}{(d+1)h_{i}^{K}}\,\|b\|_{\infty,K}.$ (49)
Similarly,
$\left|\int_{K}(b\cdot\nabla\phi_{i})(b\cdot\nabla\phi_{j})\,dx\right|$ $\leq\int_{K}|b|^{2}|\nabla\phi_{i}||\nabla\phi_{j}|\,dx$ (50)
$\leq\|b\|_{\infty,K}^{2}|\nabla\phi_{i}||\nabla\phi_{j}||K|=\frac{|K|}{h_{i}^{K}h_{j}^{K}}\,\|b\|_{\infty,K}^{2},$
which gives
$\left|\int_{K}(b\cdot\nabla\phi_{i})(b\cdot\nabla\phi_{j})\,dx\right|\leq\frac{|K|}{h_{i}^{K}h_{j}^{K}}\,\|b\|_{\infty,K}^{2}.$ (51)
For $\mathbb{D}_{K}$ in (4), following [27], we define the diffusion-dependent map $\mathbb{D}_{K}^{-1/2}{\boldsymbol{x}}:K\to\tilde{K}$. Define the $\mathbb{D}_{K}$-norm by $\|{\boldsymbol{x}}\|_{\mathbb{D}_{K}}=\sqrt{{\boldsymbol{x}}^{T}\mathbb{D}_{K}{\boldsymbol{x}}}$. Then we have the following relations
@@
Denote by $\tilde{\alpha}_{ij}^{K}$ the dihedral angle between faces $\tilde{S}_{i}^{K}$ and $\tilde{S}_{j}^{K}$. Then
@@
The difficulty with the SUPG formulation (2.1) is the coupling of the diffusion tensor with the flow field $\boldsymbol{b}$} in (27) as follows:
@@
where $\mathbb{D}_{(i)}$ is the vector $i$th row of $\mathbb{D}$. An essential assumption to the following analysis is that the tensor defined by
$\mathbb{K}=\sum_{i=1}^{d}\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}\otimes\mathbb{D}_{(i)},$ (58)
is symmetric and positive definite. This highlights a significant relationship between $\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}$ and the $i$th row of $\mathbb{D}$ when using SUPG. For each element $K\in\mathcal{T}_{h}$, and denote the elementwise average of $\mathbb{K}$ as
$\mathbb{K}_{K}=\frac{1}{|K|}\int_{K}\mathbb{K}(\boldsymbol{x})d\boldsymbol{x},$ (59)
according to (24).
In the analysis below, we assume that $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ are symmetric and positive definite. This assumption imposes a structural alignment between the velocity field and the diffusion tensor. While restrictive, it enables tractable analysis of DMP preservation. This assumption is not expected to hold in general flows but serves as a sufficient condition.
The added diffusion integral term in (2.1) can be written in general form as
$\tau_{K}\int_{K}(\mathbb{D}\nabla u)\cdot\nabla({\boldsymbol{b}}\cdot\nabla v)\textrm{d}{\boldsymbol{x}}$ $=\tau_{K}\sum_{i=1}^{d}\int_{K}(\nabla v)^{T}\left(\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}\otimes\mathbb{D}_{(i)}\right)\nabla u\textrm{d}{\boldsymbol{x}}=\tau_{K}(\nabla v)^{T}\mathbb{K}_{K}\nabla u.$ (60)
@@
$(\nabla\varphi_{i})^{T}\mathbb{F}_{K}\nabla\varphi_{j}=\begin{cases}\displaystyle-\frac{\cos(\tilde{\alpha}_{ij}^{K})}{\tilde{h}_{i}^{K}\tilde{h}_{j}^{K}}\textrm{\ \ for \ \ }i\neq j\\ \displaystyle\frac{1}{(\tilde{h}_{i}^{K})^{2}}\textrm{\ \ for \ \ i = j}\end{cases}.$ (63)
@@ 4.1 General flow forms
$\frac{\partial{\boldsymbol{b}}}{\partial x_{1}}\otimes\mathbb{D}_{(1)}=\begin{pmatrix}\displaystyle\frac{\partial b_{1}}{\partial x_{1}}D_{11}&\displaystyle\frac{\partial b_{1}}{\partial x_{1}}D_{12}\\ \displaystyle\frac{\partial b_{2}}{\partial x_{1}}D_{11}&\displaystyle\frac{\partial b_{2}}{\partial x_{1}}D_{12}\end{pmatrix}$, (65)
@@
$\frac{\partial b_{1}}{\partial x_{1}}$ $=\left(\frac{D_{11}}{D_{12}}\right)\frac{\partial b_{2}}{\partial x_{1}},$ (73)
@@
$\frac{\partial b_{1}}{\partial x_{2}}$ $=\left(\frac{D_{21}}{D_{22}}\right)\frac{\partial b_{2}}{\partial x_{2}},$ (75)
@@
$\left(\frac{D_{11}}{D_{12}}\right)b_{2}(x_{1},x_{2},t)+f_{2}(x_{2},t)$ $=\left(\frac{D_{21}}{D_{22}}\right)b_{2}(x_{1},x_{2},t)+f_{1}(x_{1},t),$ (77)
$\left(\frac{D_{11}}{D_{12}}-\frac{D_{21}}{D_{22}}\right)b_{2}(x_{1},x_{2},t)$ $=f_{1}(x_{2},t)-f_{2}(x_{1},t),$ (78)
@@
$b_{1}(x_{1},x_{2},t)$ $=\left(\frac{D_{11}}{D_{12}}\right)\frac{D_{12}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))}{\det(\mathbb{D})}+f_{2}(x_{2},t)$
$=\frac{D_{11}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))+f_{2}(x_{2},t)(D_{11}D_{22}-D_{12}D_{21})}{\det(\mathbb{D})}$
@@
$\frac{1}{\det(\mathbb{D})}\left(D_{22}f_{1}(x_{1},t)\mathbb{D}_{(1)}-D_{12}f_{2}(x_{2},t)\mathbb{D}_{(2)}\right),$ (81)
@@
$\frac{1}{\det(\mathbb{D})}\left(D_{11}g_{2}(x_{2},t)\mathbb{D}_{(2)}-D_{21}g_{1}(x_{1},t)\mathbb{D}_{(1)}\right).$ (82)
@@
${\boldsymbol{b}}(x_{1},x_{2},t)=\begin{pmatrix}\displaystyle\frac{f(x_{1},t)}{D_{11}}\\ \displaystyle\frac{g(x_{2},t)}{D_{22}}\end{pmatrix}$, (83)
@@
$\frac{D_{22}}{\det(\mathbb{D})}\left(D_{11}\frac{\partial f_{1}}{\partial x_{1}}-D_{12}\frac{\partial f_{2}}{\partial x_{2}}\right)$ (85)